A car starts from rest and covers a distance s under accelerated motion, after which the body moves with a constant velocity and a covers a distance 2s and finally comes under retarded motion and covers distance 5s and then comes to rest. Find out the ratio of Average velocity and Maximum velocity?

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
t1 + t2 + t3 = 2s/v + 2s/v + 10s/v = 14 s/v
 
Now avg velocity = 8s /(t1 + t2+ t3) = 8s/14s/v = 4v/7
 
Hence
 
Avg /maximum velocity = 4/7